OpenStudy (anonymous):

a recipe for desert calls for 2/4 cup of powered sugar and 1/6 cup of brown sugar .what is the total amount of sugar needed for the recipe

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

6/12 + 2/12 = 8/12 = 2/3

OpenStudy (anonymous):

2/4 = 3/6 + 1/6 = 4/6 = 2/3

OpenStudy (aroub):

2/4+1/6 2/4*6 +1/6*4 12/24 +4/24 16/24 8/12 4/6 2/3 hey, do u want to learn how do to do this...???

OpenStudy (aroub):

lol sorry "do u want to learn how to do this?"

OpenStudy (anonymous):

yea